5y+7-7y is less than or equal to 13
Answer by checkley77(12844)

5y+7-7y<=13
-2Y<=13-7
-2Y<=6
Y>=6/-2 NOTE THE < SIGN CHANGES TO A > SIGN WHENEVER YOU DIVIDE BY A NEGATIVE VALUE.
Y>=-3 ANS.
PROOF:
LET Y=-2.9
5*-2.9+7-7*-2.9<=13
-14.5+7+20.3<=13
12.8<=13